# Multi-color XFEL pulses with variable color separation and time delay for multi-frame diffraction imaging

### v1

@@ 3 Start-to-end XFEL simulation
Under fully optimized $R_{56}$ values for all dispersive chicanes, the undulator modules $A_{1}-A_{5}$, $B_{1}-B_{6}$, $C_{1}-C_{6}$ and $D_{1}-D_{6}$ have $K$ values of $K_{1}$ = 2.00, $K_{2}$ = 2.21, $K_{3}$ = 2.39 and $K_{4}$ = 2.56, respectively, as illustrated in Fig. 5(c). The electron beam propagating through these undulator modules generates radiation at the resonant wavelengths of 1.8, 2.1, 2.4, and 2.7 nm, with adjacent wavelengths separated by 0.3 nm, shown in Fig. 5(a). These wavelengths correspond to photon energies of 688.80, 590.40, 516.60, and 459.20 eV, respectively. Independent generation of each X-ray pulse is enabled by their time separation via delay chicanes, resulting in inter-pulse temporal separations of delay = 1 ps, evident in the FEL power profile (see Fig. 5(b)). All four-color X-ray pulses achieve peak power levels on the order of hundreds of megawatts. To realize this multi-color output within a single pass, the same electron bunch is reused across multiple undulator modules. However, the FEL radiation in the upstream sections induces significant energy spread in the high-current core of the bunch, thereby degrading its FEL gain for downstream wavelengths. Consequently, lasing at subsequent colors is sustained primarily by the low-energy-spread electrons in the bunch head and tail. This longitudinal partitioning of lasing regions leads to temporal pulse splitting in the output. As demonstrated in the following sections, a precisely balanced pulse energy configuration across colors effectively suppresses this splitting phenomenon, thereby improving temporal structure and synchronization.
Table 1: Main electron beam parameters of FEL-II
@@
Normalized emittance 0.15 $m\cdotmrad$
Slice energy spread 0.01 %
Bunch charge 100 pC
Undulator
@@
Figure 5: (a) Undulator parameter setting in the whole beamline. (b) The spectra of the multi-color XFEL pulses at the undulator module exit. (c) The various FEL power profiles at the undulator module exit with a time delay of 1 ps.
@@
Figure 6(a) (solid line) shows that the bunching factor in the first undulator module $A_{1}$ remains nearly zero, indicating the absence of significant microbunching structure in the electron beam. When optimal dispersive strengths $R_{56}^{1}=1.84\ \mu m$, $R_{56}^{2}=1.85\ \mu m$, $R_{56}^{3}=1.74\ \mu m$, and $R_{56}^{4}=1.15\ \mu m$ are implemented during the progression through the module $A_{5}$, the bunching factor undergoes four sharp vertical transitions. These transitions result in a dramatic increase in the bunching factor from the shot-noise level to approximately 0.54. These distinct jumps, under the four-stage OK configuration, signify a substantial enhancement in microbunching. The phenomenon highlights the effectiveness of using four optimized parameters $R_{56}$ to enhance energy modulation and convert it into density modulation four times within the undulator modules $A_{1}-A_{5}$. Consequently, this process significantly enhances the FEL gain of the radiation at $\lambda_{1}$ = 1.8 nm. Between undulator modules $A_{5}$ and $B_{1}$, the bunching factor gradually decays and stabilizes at lower values due to microbunching disruption by the delay chicane 1 (delay = 1 ps). Subsequent modules operating in the five-stage OK mode ($B_{1}-B_{6}$, $C_{1}-C_{6}$, $D_{1}-D_{6}$) display identical behavior: each vertical transition represents a phase of microbunching enhancement. Notably, the microbunching is disrupted by delay chicanes 2 and 3, as observed between $B_{6}$ and $C_{1}$, as well as between $C_{6}$ and $D_{1}$.
@@
Figure 6: Evolution of the maximum bunching factor (a), mean energy spread (b), and FEL pulse energy (c) along the beamline (z-direction), comparing two configurations: (1) Without OK (dot-dashed plot); (2) With OK (solid plot, plain SASE) an d based on three-color X-ray pulses that include OK, with the fourth pulse employing SASE mode (gray dashed plot).
@@
As the beam energy spread increases along the undulator modules in Fig. 6(b), progressively violating the initial condition $\sigma_{\delta}\ll\rho$. This degradation results in declining pulse energies for the different X-ray wavelengths: 30.4 (purple), 25.6 (blue), 1.9 (green), and 2.9 $\mu J$ (red). Comparative analysis with and without OK (Fig. 6) demonstrates the benefits of implementing the multi-stage OK-SASE enhancement mode in the first, second, and third pulses. For the fourth pulse, the case with OK does not yield significant FEL improvement due to its larger initial energy spread, compared to the scenario without OK (dot-dashed line), which exhibits a more gradual energy spread growth and thus a relatively smaller initial energy spread. However, when the first three pulses are enhanced by OK while the fourth pulse reverts to normal SASE (gray dashed line), the output performance deteriorates significantly. This indicates that the OK mechanism still plays a crucial role in accelerating the FEL gain process across different wavelengths. As a result, high-power four-color XFEL pulses can be generated within shorter undulator modules.
@@
4 Multi-frame diffraction imaging
Due to the fully optimized $R_{56}$ values in previous simulation, the first two colors were very strong, which in turn led to quite weak pulses for the last two colors. In order to achieve a more evenly tailored distribution of pulse energies, further fine-tuning of the appropriate $R_{56}$ of the dispersive chicanes was performed in the undulator modules. As a result, the pulse energies for the four-color X-ray pulses were successfully tailored to 11.3, 8.3, 9.0, and 9.2 $\mu J$, as shown in Fig. 8, respectively. The key FEL properties are summarized in Table 2, including the pulse energy, the full width at half maximum (FWHM) of pulse length, spectral bandwidth (FWHM), transverse beam size (rms), transverse divergence (rms). Figure 9 displays the power density distribution of the four-color X-ray pulses, from which the FWHM transverse beam size can be obtained.
Figure 8: Spectra (a) and peak power profiles (b) at the exit of the undulator module of various colors. (c) Gain curves of four-color FEL pulses.
Table 2: FEL performance of the various colors.
@@ 4 Multi-frame diffraction imaging
Figure 9: Transverse power density distributions of the four-color XFEL pulses at the undulator module exit: (a) First pulse; (b) Second pulse; (c) Third pulse; (d) Fourth pulse.
@@
To evaluate the applicability of this type of light for scattering experiments, we can assess its spectroscopic utility to some extent by analyzing its spectral distribution. In this analysis, dynamical processes within the sample and the temporal characteristics of multi-color XFEL pulses were not taken into account. For this purpose, we introduce an X-ray grating. We employ a grating with 2400 lines per millimeter, directing radiation toward the grating surface at a grazing incidence angle of 5 degrees. Here, we consider only the first diffraction order of the grating and position a detector at the image plane corresponding to the central wavelength of the FEL radiation to examine the spectral characteristics of the radiation.
@@
The grating-object distance is set to 5 m, and the image distance is 2 m. Given that our radiation comprises four distinct wavelengths — 1.8, 2.1, 2.4, and 2.7 nm — the distances traveled by each wavelength to reach the grating are 97.07, 66.38, 35.69, and 5 m. Therefore, the footprint of radiation incident on the grating surface corresponds to each wavelength is 1.6 mm for $\lambda_{1}$, 1.2 mm for $\lambda_{2}$, 0.7 mm for $\lambda_{3}$, 0.1 mm for $\lambda_{4}$. The corresponding grating resolution $R$ values were determined to be 3840, 2880, 1680, and 240, with minimum wavelength resolution $\Delta\lambda_{min}$ equal to 0.47, 0.73, 1.4, and 11.3 pm, respectively. The theoretical resolution limit of the grating for each wavelength is significantly smaller than the full width at half maximum (FWHM) bandwidth of each wavelength. Under ideal conditions, the grating should theoretically be capable of resolving four-color XFEL pulses with wavelength differences of 0.3 nm.
@@
The Shadow 4 51 optical tracing code framework was employed for optical tracing simulations of the aforementioned grating system. This framework is commonly used in the field of synchrotron radiation optics. Consequently, the relative diffraction angle differences from central wavelength ($\lambda_{0}$ = 2.25 nm) are $0.471^{\circ}$, $0.154^{\circ}$, $0.151^{\circ}$, $0.444^{\circ}$. The distances of these wavelengths from $\lambda_{0}$ are 16.45, 5.37, 5.27, and 15.52 mm, respectively, as shown in Fig. 10. These results could be recorded using a large-area X-ray CCD detector52 ($4096\times 4096$ pixels, pixel size $15\times 15\ \mu m^{2}$). Thus, this simulation process of X-ray transport design validated the feasibility of the multi-color X-ray pulse diffraction imaging scheme.
@@
Figure 10: Distinguishable diffraction images were recorded, centered around $\lambda_{0}$, with wavelengths of 1.8, 2.1, 2.4, and 2.7 nm arranged from bottom to top.
5 Conclusion and discussion
A novel multi-frame X-ray diffraction imaging method is proposed by theoretical analysis and start-to-end simulation. This approach utilizes the multi-stage OK-SASE configurations to generate high-power, multi-color XFEL pulses with tunable time delay and wavelength separation. It is further demonstrated with Shadow 4 for X-ray transport to validate the feasibility of the diffraction imaging process. The time interval between four-color X-ray pulses, controlled through delay chicanes, enables higher temporal resolution imaging. The 5 m spacing between undulator modules of different colors is sufficient to provide hundreds of femtosecond delay, even up to the picosecond level. In this discussion, generating the wavelengths of interest (1.8, 2.1, 2.4, and 2.7 nm) with a flexible wavelength separation, e.g., 0.3 nm plays a crucial role in optimizing the contrast and spatial resolution of diffraction imaging results. Furthermore, it is particularly advantageous for investigating specific elemental absorption edges, such as the oxygen K-edge at 529.0 eV. Extending the approach to other soft X-ray regions, such as 6 nm, which encompasses the carbon K-edge at 284 eV, further broadens its applicability. The precise adjustment of $R_{56}$ of the dispersive chicanes effectively controls the pulse energy of each color. This ensures a sufficient photon flux for an improved signal-to-noise ratio and imaging clarity, while simultaneously mitigating sample damage, making it particularly suitable for studying fragile specimens. Based on the performance characteristics and actual requirements of the light source, the structure of the undulator can be optimized accordingly.
In this method, reusing the same electron bunch constrains independent control over the temporal profile of individual pulses, potentially leading to pulse splitting in downstream undulator modules. This issue can be alleviated either by reducing the intensity of upstream radiation while gradually enhancing that of downstream radiation, or by increasing the wavelength separation between adjacent X-ray pulses. In this study, the residual beam energy chirp about $0.22\%$ remains smaller than the FEL gain bandwidth of $0.34\%$, thus no significant spectral broadening is observed. In addition, implementing a de-chirper enables precise manipulation of the electron beam’s longitudinal phase space, providing further flexible control over pulse characteristics.
@@ 5 Conclusion and discussion
In summary, this method employs the whole electron beam combined with split undulators in OK-SASE mode to generate four-color XFEL pulses with adjustable time delay and wavelength separation, which can be used to validate diffraction images. Additionally, the multi-stage OK configurations can be applied to other multi-color schemes, speeding up the saturation of each color. This approach is expected to provide a better foundation for designing sample-specific light sources tailored to dynamic imaging experiments for XFEL users. Nevertheless, multi-color X-ray pulses could enable applications in multi-color pump-probe experiments, nonlinear X-ray spectroscopy, multi-wavelength anomalous diffraction, and other multi-wavelength techniques. Collectively, these capabilities provide a robust framework for exploring microscopic mechanisms in ultrafast science.
Acknowledgements
@@
We extend our gratitude to Eduard Prat, Sven Reiche, Manuel Sanchez del Rio, Zipeng Liu, and Zheng Qi for their insightful discussions and keen interest in this work. This work was supported by the Natural Science Foundation of China (Grant Nos. 12175058, 12125508, 12575251), National Science Fund of Hunan Province for Distinguished Young Scholars No. 2024JJ2009, Beijing Distinguished Young Scientist Program and National Grand Instrument Project (Grant No. SQ2019YFF01014400), the National Key Research and Development Program of China (2024YFA1612101, 2024YFA1612104), Shanghai Pilot Program for Basic Research – Chinese Academy of Sciences, Shanghai Branch (JCYJ-SHFY-2021-010), and China Postdoctoral Science Foundation (2025M770914).

### v2

@@ III Start-to-end XFEL simulation
Table 1: Main electron beam and undulator parameters of FEL-II.
@@
Slice energy spread 0.01 %
Normalized emittance 0.15 $m\cdotmrad$
Bunch charge 100 pC
@@
Undulator line
@@
Under fully optimized $R_{56}$ values for all dispersive chicanes, the undulator modules $A_{5}$, $B_{6}$, $C_{6}$, $D_{6}$ (with five sub-undulators in $A_{5}$ and six sub-undulators in each of $B_{6}$, $C_{6}$ and $D_{6}$) are operated with undulator parameters $K_{1}$ = 2.00, $K_{2}$ = 2.21, $K_{3}$ = 2.39 and $K_{4}$ = 2.56, as illustrated in Fig. 5(c). The electron beam propagating through these undulator modules generates radiation at the resonant wavelengths of 1.8, 2.1, 2.4, and 2.7 nm, with adjacent wavelengths separated by 0.3 nm, shown in Fig. 5(a). These wavelengths correspond to photon energies of 688.80, 590.40, 516.60, and 459.20 eV, respectively. Independent generation of each X-ray pulse is enabled by their time separation via delay chicanes, resulting in inter-pulse temporal separations of delay = 1 ps, evident in the FEL power profile (see Fig. 5(b)). All four-color X-ray pulses achieve peak power levels on the order of hundreds of megawatts. To realize this multi-color output within a single pass, the same electron bunch is reused across multiple undulator modules. However, the FEL radiation in the upstream sections induces significant energy spread in the high-current core of the bunch, thereby degrading its FEL gain for downstream wavelengths. Consequently, lasing at subsequent colors is sustained primarily by the low-energy-spread electrons in the bunch head and tail. This longitudinal partitioning of lasing regions leads to temporal pulse splitting in the output. As demonstrated in the following sections, a precisely balanced pulse energy configuration across colors effectively suppresses this splitting phenomenon, thereby improving temporal structure and synchronization.
@@
Figure 5: (a) The spectra of the multi-color XFEL pulses at the undulator module exit. (b) The various FEL power profiles at the undulator module exit with a time delay of 1 ps. (c) Undulator parameter setting in the whole beamline.
@@
Figure 6(a) (solid line) shows that the bunching factor in the first sub-undulator $U_{A,1}$ remains nearly zero, indicating the absence of significant microbunching structure in the electron beam. When optimal dispersive strengths $R_{56}^{1}=1.84\ \mu m$, $R_{56}^{2}=1.85\ \mu m$, $R_{56}^{3}=1.74\ \mu m$, and $R_{56}^{4}=1.15\ \mu m$ are implemented during the progression through the sub-undulator $U_{A,5}$, the bunching factor undergoes four sharp vertical transitions. These transitions result in a dramatic increase in the bunching factor from the shot-noise level to approximately 0.54. These distinct jumps, under the four-stage OK configuration, signify a substantial enhancement in microbunching. The phenomenon highlights the effectiveness of using four optimized parameters $R_{56}$ to enhance energy modulation and convert it into density modulation four times within the undulator module $A_{5}$. Consequently, this process significantly enhances the FEL gain of the radiation at $\lambda_{1}$ = 1.8 nm. Between sub-undulators $U_{A,5}$ and $U_{B,1}$, the bunching factor gradually decays and stabilizes at lower values due to microbunching disruption by the delay chicane 1 (delay = 1 ps). Subsequent modules operating in the five-stage OK mode ($B_{6}$, $C_{6}$, $D_{6}$) display identical behavior: each vertical transition represents a phase of microbunching enhancement. Notably, the microbunching is disrupted by delay chicanes 2 and 3, as observed between $U_{B,6}$ and $U_{C,1}$, as well as between $U_{C,6}$ and $U_{D,1}$.
@@
Figure 6: Evolution of the slice-averaged bunching factor (a), slice-averaged energy spread (b), and FEL pulse energy (c) along the beamline (z-direction), comparing two configurations: (1) Without OK (dot-dashed plot, plain SASE); (2) With OK (solid plot) and based on three-color X-ray pulses that include OK, with the fourth pulse employing SASE mode (gray dashed plot).
@@
As the beam energy spread increases along the undulator modules in Fig. 6(b), the condition $\sigma_{\delta}\ll\rho$, to which the OK enhancement is highly sensitive [4, 14, 38], is progressively violated, leading to declining pulse energies for the different X-ray wavelengths: 30.4 $\mu J$ (1.8 nm), 25.6 $\mu J$ (2.1 nm), 1.9 $\mu J$ (2.4 nm), and 2.9 $\mu J$ (2.7 nm). In contrast, when OK chicanes are removed, the corresponding pulse energies drop to 0.2, 4.5, 1.7, and 5.5 $\mu J$. The multi-stage OK enhancement is most effective for the first three pulses. For the fourth pulse, the case with OK does not yield significant FEL improvement due to its larger initial energy spread, compared to the scenario without OK (dot-dashed line), which exhibits a more gradual energy spread growth and thus a relatively smaller initial energy spread. However, when the first three pulses are enhanced by OK while the fourth pulse reverts to normal SASE (gray dashed line), the output performance deteriorates significantly. This indicates that the OK mechanism still plays a crucial role in accelerating the FEL gain process across different wavelengths. As a result, high-power four-color XFEL pulses can be generated within shorter undulator modules.
@@
IV Multi-frame diffraction imaging
It is necessary to characterize the statistical properties and main parameters of the generated four-color XFEL pulses under realistic operating conditions. In particular, the pulse-energy balance among different colors, the shot-to-shot stability inherent to the SASE process, and the transverse radiation properties play a crucial role in determining the applicability of the source for downstream experiments.
Since the $R_{56}$ values were fully optimized in the previous simulation, the first two colors exhibited substantially higher pulse energies, at the expense of the last two colors, which became relatively weak. To obtain a more evenly tailored pulse-energy distribution among the four colors, additional fine-tuning of the $R_{56}$ values of the dispersive chicanes in the undulator modules was performed. To account for shot-to-shot fluctuations inherent to SASE, the pulse energies were evaluated based on 50-shot simulations. The results, shown in Fig. 8(a), demonstrate the overall stability of the FEL performance, whereas Figs. 8(b) and (c) present a representative single shot illustrating the typical pulse characteristics. Figure 9 displays the corresponding transverse power density distributions of the four-color X-ray pulses, from which the rms transverse spot sizes are extracted. The main FEL properties of this representative shot are summarized in Table 2, including pulse energy, pulse length (FWHM), spectral bandwidth (FWHM), transverse spot size (rms), and transverse divergence (rms). These parameters serve as the input conditions for the subsequent X-ray transport simulations.
Figure 8: (a) Pulse energies of the four XFEL pulses from 50-shot simulations. (b) Spectra and (c) power profiles of the four-color XFEL pulses at the exit of the undulator modules for a representative single shot.
Table 2: FEL performance parameters for pulses at different wavelengths.
@@ IV Multi-frame diffraction imaging
Figure 9: Transverse intensity profiles of the four-color XFEL pulses at the exit of the undulator module for a representative shot: (a) first pulse, (b) second pulse, (c) third pulse, and (d) fourth pulse.
@@
To assess the applicability of this radiation for scattering experiments, its spectroscopic performance is further investigated by analyzing the spectral distribution. In this analysis, neither the dynamical response of the sample nor the temporal structure of the multi-color XFEL pulses is considered, allowing the focus to be placed solely on the intrinsic spectral properties of the radiation. An X-ray grating with 2400 lines/mm is employed, with the beam incident on the grating surface at a grazing angle of $5^{\circ}$. Only the first diffraction order is taken into account, and a detector is placed at the image plane corresponding to the central wavelength of the FEL radiation to characterize the resulting spectral distribution.
@@
The grating-object distance is set to 5 m, and the image distance is 2 m. Given that our radiation comprises four distinct wavelengths — 1.8, 2.1, 2.4, and 2.7 nm — the distances traveled by each wavelength to reach the grating are 97.07, 66.38, 35.69, and 5 m. Therefore, the footprint of radiation incident on the grating surface corresponds to each wavelength is 1.6 mm for $\lambda_{1}$, 1.2 mm for $\lambda_{2}$, 0.7 mm for $\lambda_{3}$, 0.1 mm for $\lambda_{4}$. The corresponding grating resolution $R$ values were determined to be 3840, 2880, 1680, and 240, with minimum wavelength resolution $\Delta\lambda_{min}$ equal to 0.47, 0.73, 1.4, and 11.3 pm, respectively. The theoretical resolution limit of the grating for each wavelength is significantly smaller than the FWHM bandwidth of each wavelength. Under ideal conditions, the grating should theoretically be capable of resolving four-color XFEL pulses with wavelength differences of 0.3 nm.
@@
The Shadow 4 [12] optical tracing code framework was employed for optical tracing simulations of the aforementioned grating system. This framework is commonly used in the field of synchrotron radiation optics. Consequently, the relative diffraction angle differences from central wavelength ($\lambda_{0}$ = 2.25 nm) are $0.471^{\circ}$, $0.154^{\circ}$, $0.151^{\circ}$, $0.444^{\circ}$. The distances of these wavelengths from $\lambda_{0}$ are 16.45, 5.37, 5.27, and 15.52 mm, respectively, as shown in Fig. 10. These results could be recorded using a large-area X-ray CCD detector[17] ($4096\times 4096$ pixels, pixel size $15\times 15\ \mu m^{2}$). Thus, this simulation process of X-ray transport design validated the feasibility of the multi-color X-ray pulse diffraction imaging scheme. More importantly, realizing such single-shot, multi-frame probing within the SHINE relies on the introduction of a multi-stage OK configuration, which substantially reduces the effective undulator length required for each color. If repeated pump–probe delay scans are needed, the high repetition rate of SHINE can support this requirement by providing sufficient photon statistics.
@@
Figure 10: Distinguishable diffraction images were clearly resolved in the simulated image plane, centered around $\lambda_{0}$, with wavelengths of 1.8, 2.1, 2.4, and 2.7 nm arranged from bottom to top.
V Conclusion and discussion
A novel single-shot multi-frame X-ray diffraction imaging scheme is proposed by theoretical analysis and start-to-end simulation. The scheme employs the whole electron bunch in a multi-stage OK-SASE configuration with split undulators to generate high-power, four-color XFEL pulses featuring independently tunable wavelength separations and controllable inter-pulse delays. Delay chicanes placed in the gaps between undulator modules operating at different resonant wavelengths enable time delays on the order of one hundred femtoseconds, extendable into the picosecond regime. The feasibility of the imaging process is confirmed by X-ray transport simulations, which demonstrate that the resulting diffraction signals from distinct probe pulses can be spatially resolved and recorded simultaneously, thereby validating the scheme’s capability for capturing ultrafast dynamics in a single shot.
Furthermore, the capability to flexibly generate wavelengths of interest, such as 1.8, 2.1, 2.4, and 2.7 nm, with a tunable separation (e.g., 0.3 nm) plays an important role in optimizing both the contrast and spatial resolution of diffraction imaging. This flexibility is particularly advantageous for probing elemental absorption edges, for example, the oxygen K-edge at 529 eV. Moreover, extending the scheme to other soft X-ray regions, such as around 6 nm covering the carbon K-edge at 284 eV, further broadens its applicability to a wide range of scientific problems. Moreover, the pulse energy of each color can be effectively controlled through precise adjustment of the $R_{56}$ values of the dispersive chicanes. This enables the delivery of sufficient photon flux to improve the signal-to-noise ratio and imaging fidelity, while simultaneously mitigating radiation damage to samples, which is especially beneficial for studies of fragile specimens. Since all pulses are generated from a single electron bunch, independent control of their temporal profiles is inherently limited, and residual energy modulation may induce pulse splitting in subsequent undulator modules. This effect may be mitigated by reducing the radiation intensity in upstream stages while gradually increasing it in downstream stages, or by enlarging the wavelength separation between adjacent pulses. In addition, the residual beam energy chirp of about 0.22% remains smaller than the FEL gain bandwidth of 0.34%, so no significant spectral broadening is observed. The implementation of a dechirper would allow more precise manipulation of the longitudinal phase space of the electron beam, offering further flexibility in tailoring pulse characteristics.
@@ V Conclusion and discussion
The multi-stage OK configuration can be extended to other multi-color approaches, such as fresh-slice operation, where it can further enhance FEL gain and enable higher peak power or shorter pulses. In particular, under low-charge operation, where fresh-slice manipulation becomes challenging, the full-beam scheme combined with OK may still provide a practical route to two-color FEL generation. Nevertheless, the proposed scheme is expected to provide a better foundation for designing sample-specific light sources tailored to dynamic imaging experiments for XFEL users, with its multi-color X-ray pulses enabling applications such as multi-color pump-probe experiments, nonlinear X-ray spectroscopy, multi-wavelength anomalous diffraction, and other multi-wavelength techniques. Collectively, these capabilities provide a robust framework for exploring microscopic mechanisms in ultrafast science.
Acknowledgements.
@@
We extend our gratitude to Eduard Prat, Sven Reiche, Manuel Sanchez del Rio, Zipeng Liu, and Zheng Qi for their insightful discussions and keen interest in this work. J. Q. Y. would like to thank the support from the junior visiting professor program of Peng Huanwu Innovation Research Center for Theoretical Physics in 2025. This work was supported by the Natural Science Foundation of China (Grant Nos. 12175058, 12125508, 12541503, 12575251), National Science Fund of Hunan Province for Distinguished Young Scholars No. 2024JJ2009, the National Key Research and Development Program of China (2024YFA1612104), Beijing Distinguished Young Scientist Program and National Grand Instrument Project (Grant No. SQ2019YFF01014400), Shanghai Pilot Program for Basic Research – Chinese Academy of Sciences, Shanghai Branch (JCYJ-SHFY-2021-010), and China Postdoctoral Science Foundation (2025M770914).